# Quantum Rainbow Codes

### v2

@@ V.1 Figure-eight graphs
which is equivalent to a pair of length-4 cycles glued at a single level-$1$ vertex, and so by ungluing these cycles and taking the Cartesian product we can obtain a flag graph equivalent to eight disjoint 3D colour code lattices on 3-torii, each of which have parameters $[\![384,9,4]\!]$. The flag grap of $\mathcal{G}_{\square}$ is then equivalent to the joining of these eight lattices at splittable seams as in Fig. 10, which also shows some example logical operator structures.
@@
We can count the number of physical qubits in this case as it is simply eight times the number of qubits in a single one of the 3D colour codes, so $n=3072$. The number of logical qubits is $k=24$ by Eq. 26 (as all $n_{c}^{i}=2$) and the distance is twice the distance of any of the component colour codes so $d=8$, so we have a $[\![3072,24,8]\!]$ code. In contrast, the 3D colour code defined from a product of three length-8 cycle graphs is a $[\![3072,9,8]\!]$ code, although this is not the most efficient colour code for this $k$ and $d$ as there is also a $[\![768,9,8]\!]$ code that is not directly obtained from the hypergraph product (we show in the Appendix A how to obtain this code from the $[\![3072,9,8]\!]$ code).
We can also try to understand the action of transversal $T/T^{\dagger}$ on this code. With respect to the partitioning of the cube shown in Fig. 10 we can label each component colour code as being either front or back (F or B), left or right (L or R), and up or down (U or D). Logical operators of colour $c_{1}$ or $c_{2}$ in the rainbow code can be described by a single one of these labels while logical operators of colour $c_{0}$ can be described by a pair (e.g. the logicals shown in Fig. 10 are the $c_{1}$, $c_{2}$ and $c_{0}$ operators). We can use this to draw an interaction graph between the logical qubits describing the logical action of transversal $T/T^{\dagger}$ (Fig. 11). We can see from this graph that two logical $CCZ$ gates are applied to each $c_{0}$ logical qubit while four $CCZ$ gates are applied to each $c_{1}$ and $c_{2}$ logical qubit. This example is available in the linked Jupyter notebook.
@@ V.2 Fully connected bipartite graphs
whereas the size of a fundamental cycle basis (by Eq. 21) is $n_{c}=16-8+1=9$. By Eq. 26 we then have 297 logical qubits in a product of three of these codes. To count the number of physical qubits we can start by noticing that, for a product of $D$ $d$-regular graphs, the number of flags associated to each level-00 node in the product graph is
@@
This is because the first (level-00) vertex in each flag is $(g_{1,i}^{0},g_{2,i}^{0},...,g_{D,i}^{0})^{0}$ and then, to find a connected level-$1$ vertex, we need to change one of the $g_{j,i}^{0}$ to $g_{j,i^{\prime}}^{1}$. There are $D$ choices of $j$ and then $d$ choices of vertex for each $j$, each with form $((g_{1,i}^{0},...g_{j,i}^{1},...,g_{D,i}^{0})^{1}$. For the level-$2$ vertex we have $(D-1)$ choices of $j$ and $d$ choices of vertex for each $j$ and so on, eventually giving $D!d^{D}$. The total number of level-00 nodes in the product graph is
@@
where $n_{0}^{i}$ is the number of level-00 nodes in input graph $i$. The total number flags is then
@@
and in this case we have $D=3,d=n_{0}^{i}=4$ so $n=4^{3}\times 3!\times 4^{3}=24576$. The distance of this code, as with all mixed HGP codes, is twice the minimum girth of any input graph, which is $2\times 4=8$, so we have a code with parameters $[\![24576,297,8]\!]$. To achieve the same $k$ and $d$ with 3D colour codes we would need $33$ copies of the $[\![768,9,8]\!]$ code, which would require $25344$ physical qubits. This example is available in the linked Jupyter notebook.
@@ V.3 Expander graphs
Next we consider an asymptotic construction based on $d$-regular bipartite expander graphs for even $d$, such as those presented in lubotzky_ramanujan_1988 . Hypergraph product codes defined using such graphs can have fairly good parameters and so we might expect the corresponding rainbow codes to also scale well. Unfortunately, the parameters of these codes actually turn out to be quite bad (worse than those of a “code” corresponding to $\Theta(n)$ copies of the Euclidean 3D colour code). Specifically, as $n_{0}^{i}$ is linear in the size $n$ of the input graphs the a code obtained from the product of three of these graphs will have $\Theta(n^{3})$ physical qubits. The number of independent cycles in these graphs is also linear in $n$ as the number of edges in a $d$-regular graph is $nd/2$, and so by Eq. 26 the number of logical qubits is $\Theta(n^{2})$. Finally, these graphs have girth $O(\log(n))$ and so the distance has the same scaling and we have a code family with parameters $[\![n,\Theta(n^{2/3}),\Theta(\log(n)]\!]$ after rescaling $n^{3}\rightarrow n$.
@@ V.4 Expander graph and a hyperbolic cellulation
Finally we present a family of finite rate and non-constant distance codes combining HGP rainbow codes with the quasi-hyperbolic 3D colour codes of zhu2023non , which were obtained from manifolds corresponding to the product of a 2D hyperbolic manifold and a circle. This is compatible with our construction as cellulations of the hyperbolic manifold can be represented as graphs with three levels of vertex, but because these graphs are not themselves products of bipartite graphs the parameters of the resulting code will not be limited by Eq. 26. The hyperbolic manifold has area $A$, genus $g=\Theta(A)$ and systole $\Theta(\log(A))$, and so a topological code defined on this manifold has parameters $[\![n,\Theta(n),\Theta(\log(n))]\!]$. The circle has length $\Theta(\log(A))$ and so the code defined on the product manifold has parameters $[\![n,\Theta(n/\log(n)),\Theta(\log(n))]\!]$ as the product with the circle increases the number of physical qubits by a factor of $\Theta(\log(n))$ while only increasing the number of logical qubits by a constant factor. We can visualise this manifold as in Fig. 12 a), where the $x$ and $y$ directions have a hyperbolic metric and the $z$ direction has a Euclidean metric. A colour code defined on this manifold (using the same subdivision of a chain complex into flags that we have described above) supports a transversal non-Clifford gate implemented by $T/T^{\dagger}$ and whose logical action is $CCZ$ between triples of qubits with logical $Z$ operators $\{Z_{c_{i}},Z_{c_{j}},Z_{c_{k}}\}$ where $Z_{c_{i}}$ and $Z_{c_{j}}$ surround a common handle in the 2D manifold while $Z_{c_{k}}$ is the unique $z$-direction logical of colour $c_{k}\neq c_{i}\neq c_{j}$. Examples of $Z$ and $X$ operators for such a triple are also shown in Fig. 12 a). Readers who desire a more thorough explanation of these properties can consult section V A of zhu2023non .
@@
Figure 12: a) Examples of logical $Z$ (above) and $X$ (below) operators in a quasi-hyperbolic colour code as defined in zhu2023non . The circular cross section in the $(x,y)$ plane is a closed 2D hyperbolic manifold and contains an extensive number of handles, while the $z$ direction is Euclidean with top and bottom faces of the cylinder associated. The triple of $X$ logicals shown will be acted on by logical $CCZ$ when transversal $T/T^{\dagger}$ is applied to the code. b) Examples of logical operators in a product of a 2D hyperbolic manifold and an expander graph. We choose a convention red = $c_{0}$, green = $c_{1}$, blue = $c_{2}$. (Left) Nontrivial $c_{1}$ and $c_{2}$ membranes perpendicular to seams meet at a nontrivial $c_{0}$ string. (Right) a nontrivial perpendicular $c_{1}$ and parallel $c_{2}$ membrane meet at a nontrivial $c_{0}$ string. All such $c_{0}$ strings originating from different quasi-hyperbolic codes correspond to the supports of independent logical operators.
@@
To achieve linear encoding rate we can instead consider the product of a $d$-regular (for even $d$) bipartite expander graph of size $s$ and a 2D hyperbolic manifold with area $A$. A rainbow code defined on the output of this product would have a number of physical qubits $\Theta(As)$. Using the fact that the expander graph contains $\Theta(s)$ independent cycles we can view this code as a joining of $\Theta(s)$ copies of the quasi-hyperbolic code described above, and prior to being joined these copies encode $\Theta(As)$ logical qubits. This product then results in a code with linear rate as long as only a constant fraction of logical qubits are lost in the joining. We can see that this is true when using the mixed stabiliser assignment as distinct $c_{0}$-coloured logical $Z$ operators parallel to the seams (i.e. in the plane of the hyperbolic manifold) are not associated by the joining, and there are $\Theta(A)$ such operators in each quasi-hyperbolic code. The distance of this code will be the minimum of $\Theta(\log(A))$ and $\Theta(\log(s))$ so we can obtain the best relative distance by setting $A=s$, which gives a code with parameters $[\![n,\Theta(n),\Theta(\log(n))]\!]$. Figure 12 b) shows examples of triples of logical operators of this code with intersections that allow for transformation by logical $CCZ$.
@@ VI Discussion
Although we have focused mostly on simplicial complexes obtained from the hypergraph product, there exist many other methods of generating suitable complexes, e.g. coset complexes KAUFMAN2023103696 or more sophisticated product constructions breuckmann_balanced_2021 . The potential of these other kinds of complex is exemplified by our example of a family of constant rate and growing distance codes with transversal non-Clifford gates, which were derived from a complex obtained by combining a graph product with a cellulation of a hyperbolic manifold. This construction gives the first family of quantum LDPC codes defined on qubits with linear rate, growing distance and transversal non-Clifford gates, which are necessary conditions to achieve $\gamma\rightarrow 0$. The question of whether it is possible to achieve a polynomial scaling of the distance while maintaining a constant rate remains an important open problem.
@@ Appendix A Edge contraction
The smallest 3D colour code obtainable from the HGP rainbow code construction described previously is a $[\![384,9,4]\!]$ code obtained using a product of three length-$4$ cycle graphs. However, a more efficient colour code with parameters $[\![96,9,4]\!]$ defined on a different lattice is known to exist. Is there a way to obtain this code from a product construction, and if so, can a similar approach be used to generate more efficient rainbow codes?
@@
$\{g_{i}\Leftarrow g_{j}\leavevmode\nobreak\ \forall\leavevmode\nobreak\ (g_{i},g_{j})\in P\}:(\mathcal{G}/\{c_{i}\})$ (37)
@@
In other words, $\mathrm{Cont}(c_{i})$ removes from $\mathcal{G}$ all $c_{i}$-coloured edges and the glues together all vertices which were previously part of the same $\{c_{i}\}$-maximal subgraph. Another way to understand it is as the contraction of all $c_{i}$-coloured edges (hence the name), so that all vertices previously connected by $c_{i}$ edges become associated and all $c_{i}$ edges in the graph are removed. As contractions for different $c_{i}$ only affect edges of different colours these operations necessarily commute and we can also talk more generally about $\{c_{i},c_{j},...\}$-contractions ($\mathrm{Cont}(c_{i},c_{j},...)$) without needing to specify an order. An example of $c_{0}$-contraction in a 2D colour code lattice is shown in Fig. 13.
@@
[…]
@@
Figure 14: (Left) Subgraph of the flag graph of a 3D cubic lattice, equivalent to a 3D colour code lattice. Colours are $c_{0}=$ dark red, $c_{1}=$ light blue, $c_{2}=$ dark blue, $c_{3}=$ light orange, so this subgraph contains two $\{c_{0},c_{1},c_{2}\}$-maximal ($=$ rainbow) subgraphs and a $\{c_{0},c_{1},c_{3}\}$-maximal ($=$ rainbow) subgraph. (Middle) Graph obtained from contracting all $c_{0}$ edges of the previous graph, so that the vertices at the endpoints of these edges become associated and the edges themselves are deleted from the graph. The result is a subregion of a different 3D colour code lattice (up to recolouring of edges). (Right) Graph obtained from contracting all $c_{3}$ edges of the previous graph. This is also a subregion of a different colour code lattice (up to recolouring of edges).
[…]
@@
This is not the only requirement, however. If we consider a graph $\mathcal{G}$ obtained from a product of a length-$4$ and a length-$6$ cycle which defines a $[\![48,4,4]\!]$ 2D colour code then the graph $\mathrm{Cont}(c_{0}):\mathcal{G}$ defines a $[\![24,2,4]\!]$ code, so two logical qubits have been lost. We can see this in Fig. 15. We observe similar results in 3D, for example a product of three length-$6$ cycles defines a $[\![1296,9,6]\!]$ code but $c_{0}$ contraction results in a $[\![648,6,6]\!]$ code and an additional $c_{3}$ contraction results in a $[\![324,6,6]\!]$ code. This suggests that for the number of logical qubits to be preserved under contraction we should choose input graphs whose fundamental cycle bases contain only cycles of length $0\mod 4$ and not $2\mod 4$.
[…]

### v3

@@ V.1 Figure-eight graphs
which is equivalent to a pair of length-4 cycles glued at a single level-$1$ vertex, and so by ungluing these cycles and taking the Cartesian product we can obtain a flag graph equivalent to eight disjoint 3D colour code lattices on 3-torii, each of which have parameters [[384,9,4]][\![384,9,4]\!]. The flag grap of $\mathcal{G}_{\square}$ is then equivalent to the joining of these eight lattices at splittable seams as in Fig.˜10, which also shows some example logical operator structures.
@@
We can count the number of physical qubits in this case as it is simply eight times the number of qubits in a single one of the 3D colour codes, so $n=3072$. The number of logical qubits is $k=24$ by Eq.˜26 (as all $n_{c}^{i}=2$) and the distance is twice the distance of any of the component colour codes so $d=8$, so we have a [[3072,24,8]][\![3072,24,8]\!] code. In contrast, the 3D colour code defined from a product of three length-8 cycle graphs is a [[3072,9,8]][\![3072,9,8]\!] code, although this is not the most efficient colour code for this $k$ and $d$ as there is also a [[768,9,8]][\![768,9,8]\!] code that is not directly obtained from the hypergraph product (we show in the Appendix˜A how to obtain this code from the [[3072,9,8]][\![3072,9,8]\!] code).
We can also try to understand the action of transversal $T/T^{\dagger}$ on this code. With respect to the partitioning of the cube shown in Fig.˜10 we can label each component colour code as being either front or back (F or B), left or right (L or R), and up or down (U or D). Logical operators of colour $c_{1}$ or $c_{2}$ in the rainbow code can be described by a single one of these labels while logical operators of colour $c_{0}$ can be described by a pair (e.g. the logicals shown in Fig.˜10 are the $c_{1}$F, $c_{2}$U and $c_{0}$UF operators). We can use this to draw an interaction graph between the logical qubits describing the logical action of transversal $T/T^{\dagger}$ (Fig.˜11). We can see from this graph that two logical $CCZ$ gates are applied to each $c_{0}$ logical qubit while four $CCZ$ gates are applied to each $c_{1}$ and $c_{2}$ logical qubit. This example is available in the linked Jupyter notebook.
@@ V.2 Fully connected bipartite graphs
whereas the size of a fundamental cycle basis (by Eq.˜21) is $n_{c}=16-8+1=9$. By Eq.˜26 we then have 297 logical qubits in a product of three of these codes. To count the number of physical qubits we can start by noticing that, for a product of $D$ $d$-regular graphs, the number of flags associated to each level-0 node in the product graph is
@@
This is because the first (level-0) vertex in each flag is $(g_{1,i}^{0},g_{2,i}^{0},...,g_{D,i}^{0})^{0}$ and then, to find a connected level-$1$ vertex, we need to change one of the $g_{j,i}^{0}$ to $g_{j,i^{\prime}}^{1}$. There are $D$ choices of $j$ and then $d$ choices of vertex for each $j$, each with form $((g_{1,i}^{0},...g_{j,i}^{1},...,g_{D,i}^{0})^{1}$. For the level-$2$ vertex we have (D1)(D-1) choices of $j$ and $d$ choices of vertex for each $j$ and so on, eventually giving $D!d^{D}$. The total number of level-0 nodes in the product graph is
@@
where $n_{0}^{i}$ is the number of level-0 nodes in input graph $i$. The total number flags is then
@@
and in this case we have $D=3,d=n_{0}^{i}=4$ so $n=4^{3}\times 3!\times 4^{3}=24576$. The distance of this code, as with all mixed HGP codes, is twice the minimum girth of any input graph, which is $2\times 4=8$, so we have a code with parameters [[24576,297,8]][\![24576,297,8]\!]. To achieve the same $k$ and $d$ with 3D colour codes we would need $33$ copies of the [[768,9,8]][\![768,9,8]\!] code, which would require $25344$ physical qubits. This example is available in the linked Jupyter notebook.
@@ V.3 Expander graphs
Next we consider an asymptotic construction based on $d$-regular bipartite expander graphs for even $d$, such as those presented in lubotzky_ramanujan_1988. Hypergraph product codes defined using such graphs can have fairly good parameters and so we might expect the corresponding rainbow codes to also scale well. Unfortunately, the parameters of these codes actually turn out to be quite bad (worse than those of a “code” corresponding to $\Theta(n)$ copies of the Euclidean 3D colour code). Specifically, as $n_{0}^{i}$ is linear in the size $n$ of the input graphs a code obtained from the product of three of these graphs will have $\Theta(n^{3})$ physical qubits. The number of independent cycles in these graphs is also linear in $n$ as the number of edges in a $d$-regular graph is $nd/2$, and so by Eq.˜26 the number of logical qubits is $\Theta(n^{2})$. Finally, these graphs have girth $O(\log(n))$ and so the distance has the same scaling and we have a code family with parameters $n,\Theta(n^{2/3}),\Theta(\log(n)]\!]$ after rescaling $n^{3}\rightarrow n$.
@@ V.4 Expander graph and a hyperbolic cellulation
Finally we present a family of finite rate and non-constant distance codes combining HGP rainbow codes with the quasi-hyperbolic 3D colour codes of zhu2023non, which were obtained from manifolds corresponding to the product of a 2D hyperbolic manifold and a circle. This is compatible with our construction as cellulations of the hyperbolic manifold can be represented as graphs with three levels of vertex, but because these graphs are not themselves products of bipartite graphs the parameters of the resulting code will not be limited by Eq.˜26. The hyperbolic manifold has area $A$, genus $g=\Theta(A)$ and systole $\Theta(\log(A))$, and so a topological code defined on this manifold has parameters $n,\Theta(n),\Theta(\log(n))]\!]$. The circle has length $\Theta(\log(A))$ and so the code defined on the product manifold has parameters $n,\Theta(n/\log(n)),\Theta(\log(n))]\!]$ as the product with the circle increases the number of physical qubits by a factor of $\Theta(\log(n))$ while only increasing the number of logical qubits by a constant factor. We can visualise this manifold as in Fig.˜12 a), where the $x$ and $y$ directions have a hyperbolic metric and the $z$ direction has a Euclidean metric. A colour code defined on this manifold (using the same subdivision of a chain complex into flags that we have described above) supports a transversal non-Clifford gate implemented by $T/T^{\dagger}$ and whose logical action is $CCZ$ between triples of qubits with logical $Z$ operators $\{Z_{c_{i}},Z_{c_{j}},Z_{c_{k}}\}$ where $Z_{c_{i}}$ and $Z_{c_{j}}$ surround a common handle in the 2D manifold while $Z_{c_{k}}$ is the unique $z$-direction logical of colour $c_{k}\neq c_{i}\neq c_{j}$. Examples of $Z$ and $X$ operators for such a triple are also shown in Fig.˜12 a). Readers who desire a more thorough explanation of these properties can consult section V A of zhu2023non.
@@
Figure 12: a) Examples of logical $Z$ (above) and $X$ (below) operators in a quasi-hyperbolic colour code as defined in zhu2023non. The circular cross section in the $(x,y)$ plane is a closed 2D hyperbolic manifold and contains an extensive number of handles, while the $z$ direction is Euclidean with top and bottom faces of the cylinder associated. The triple of $X$ logicals shown will be acted on by logical CCZCCZ when transversal $T/T^{\dagger}$ is applied to the code. b) Examples of logical operators in a product of a 2D hyperbolic manifold and an expander graph. We choose a convention red = $c_{0}$, green = $c_{1}$, blue = $c_{2}$. (Left) Nontrivial $c_{1}$ and $c_{2}$ membranes perpendicular to seams meet at a nontrivial $c_{0}$ string. (Right) a nontrivial perpendicular $c_{1}$ and parallel $c_{2}$ membrane meet at a nontrivial $c_{0}$ string. All such $c_{0}$ strings originating from different quasi-hyperbolic codes correspond to the supports of independent logical operators.
@@
To achieve linear encoding rate we can instead consider the product of a $d$-regular (for even $d$) bipartite expander graph of size $s$ and a 2D hyperbolic manifold with area $A$. A rainbow code defined on the output of this product would have a number of physical qubits $\Theta(As)$. Using the fact that the expander graph contains $\Theta(s)$ independent cycles we can view this code as a joining of $\Theta(s)$ copies of the quasi-hyperbolic code described above, and prior to being joined these copies encode $\Theta(As)$ logical qubits. This product then results in a code with linear rate as long as only a constant fraction of logical qubits are lost in the joining. We can see that this is true when using the mixed stabiliser assignment as distinct $c_{0}$-coloured logical $Z$ operators parallel to the seams (i.e. in the plane of the hyperbolic manifold) are not associated by the joining, and there are $\Theta(A)$ such operators in each quasi-hyperbolic code. The distance of this code will be the minimum of $\Theta(\log(A))$ and $\Theta(\log(s))$ so we can obtain the best relative distance by setting $A=s$, which gives a code with parameters $n,\Theta(n),\Theta(\log(n))]\!]$. Figure˜12 b) shows examples of triples of logical operators of this code with intersections that allow for transformation by logical $CCZ$.
@@ VI Discussion
Although we have focused mostly on simplicial complexes obtained from the hypergraph product, there exist many other methods of generating suitable complexes, e.g. coset complexes KAUFMAN2023103696 or more sophisticated product constructions breuckmann_balanced_2021. The potential of these other kinds of complex is exemplified by our example of a family of constant rate and growing distance codes with transversal non-Clifford gates, which were derived from a complex obtained by combining a graph product with a cellulation of a hyperbolic manifold. This construction gives the first family of quantum LDPC codes defined on qubits with linear rate, growing distance and transversal non-Clifford gates, which are necessary conditions to achieve $\gamma\rightarrow 0$. We also note that we have studied only a small subset of the many possible stabiliser assignments permitted by our construction, and that there may be other choices that result in codes with comparable or superior properties. The identification of more sophisticated ways to choose or understand such assignments is left as an open problem.
@@ Appendix A Edge contraction
The smallest 3D colour code obtainable from the HGP rainbow code construction described previously is a [[384,9,4]][\![384,9,4]\!] code obtained using a product of three length-$4$ cycle graphs. However, a more efficient colour code with parameters [[96,9,4]][\![96,9,4]\!] defined on a different lattice is known to exist. Is there a way to obtain this code from a product construction, and if so, can a similar approach be used to generate more efficient rainbow codes?
@@
$\{g_{i}\Leftarrow g_{j}\penalty 10000\ \forall\penalty 10000\ (g_{i},g_{j})\in P\}:(\mathcal{G}/\{c_{i}\})$ (37)
where $P$ is the set of edges of a spanning forest of $\mathcal{G}$ that contains only $c_{i}$ edges. We say that the graph $\mathrm{Cont}(c_{i}):\mathcal{G}$ has colours $\{c_{0},...,\cancel{c_{i}},...,c_{k}\}$.
@@
Contraction also defines a mapping of each maximal or rainbow subgraph via the contraction of all $c_{i}$ edges in this subgraph. For example, in Fig.˜13 the light blue/dark blue octagons are $\{c_{1},c_{2}\}$-maximal/rainbow subgraphs just as in the original code while the light blue diamonds are $\{\cancel{c_{0}},c_{1}\}$-maximal/rainbow subgraphs and the dark blue edges are $\{\cancel{c_{0}},c_{2}\}$-maximal/rainbow subgraphs. This defines a corresponding mapping on the stabilisers of the original code but in general these operators will not commute in the contracted code. For example, the intersection of a $\{\cancel{c_{0}},c_{1}\}$- and $\{\cancel{c_{0}},c_{2}\}$-maximal/rainbow subgraph in Fig.˜13 can be only a single vertex. We must therefore choose some commuting subset of these operators to define the stabilisers for a code on the contracted lattice. In the example of Fig.˜13 the most obvious choice is the operators defined on $\{\cancel{c_{0}},c_{1}\}$- and $\{c_{1},c_{2}\}$-maximal/rainbow subgraphs. We then obtain a rotated version of the original code with parameters [[16,4,4]][\![16,4,4]\!] as opposed to the [[32,4,4]][\![32,4,4]\!] uncontracted code.
The [[384,9,4]][\![384,9,4]\!] and [[96,9,4]][\![96,9,4]\!] 3D colour codes are also related by an edge contraction ($\mathrm{Cont}(c_{0},c_{3})$) as shown in Fig.˜14. The $X$ stabilisers of the contracted code are $\{\cancel{c_{0}},c_{1},c_{2}\}$- and $\{c_{1},c_{2},\cancel{c_{3}}\}$-maximal/rainbow subgraphs and the $Z$ stabilisers are $\{\cancel{c_{0}},c_{1}\}$-, $\{c_{2},\cancel{c_{3}}\}$- and $\{c_{1},c_{2}\}$-maximal/rainbow subgraphs. Because we have contracted two colours of edge we get a factor of four reduction in the number of physical qubits ($384/4=96$). There is also an intermediate lattice obtained by performing only one of the two contractions (it does not matter which one) which defines a code with parameters [[192,9,4]][\![192,9,4]\!]. We can also contract either of the remaining edge colours to obtain a octahedral-cuboctahedral lattice which supports a triple of 3D surface codes related to the 3D colour code by an unfolding map, and which admit a transversal $CCZ$ gate [vasmer_three-dimensional_2019]. The stabilisers of these three codes can also be obtained from a suitable choice of contracted subgraphs. Construction of 3D colour codes with the various edge contractions can be explored in the linked Jupyter notebook.
@@
Figure 14: (Left) Subgraph of the flag graph of a 3D cubic lattice, equivalent to a 3D colour code lattice. Colours are $c_{0}=$ dark red, $c_{1}=$ light blue, $c_{2}=$ dark blue, $c_{3}=$ light orange, so this subgraph contains two $\{c_{0},c_{1},c_{2}\}$-maximal (== rainbow) subgraphs and a $\{c_{0},c_{1},c_{3}\}$-maximal (== rainbow) subgraph. (Middle) Graph obtained from contracting all $c_{0}$ edges of the previous graph, so that the vertices at the endpoints of these edges become associated and the edges themselves are deleted from the graph. The result is a subregion of a different 3D colour code lattice (up to recolouring of edges). (Right) Graph obtained from contracting all $c_{3}$ edges of the previous graph. This is also a subregion of a different colour code lattice (up to recolouring of edges).
Figure 15: (Left) a 2D colour code defined on the flag graph of a $3\times 2$ square lattice on a torus. (Right) $c_{0}$-contraction of this lattice resulting in a code with $X$ and $Z$ stabilisers on $\{\cancel{c_{0}},c_{1}\}$- and $\{c_{1},c_{2}\}$-maximal/rainbow subgraphs. Shown in grey is a logical operator defined on pair of edges (thick light blue lines) and the stabilisers anticommuting with these edge operators (grey crosses). The operator must wrap twice around the torus in order to close, resulting in a loss of two encoded qubits relative to the uncontracted code.
In order to understand when and how this procedure might be generalised to other rainbow codes we first need to understand what makes it work in the cases discussed above. The most obvious requirement is that the stabilisers of the code must still have even weight after the edge contraction (this is necessary for the resulting code to still have a transversal non-Clifford gate due to the requirements discussed in the previous section). It is for this reason that we can contract the $c_{0}$ and $c_{3}$ edges of the [[384,9,4]][\![384,9,4]\!] code but not the $c_{1}$ or $c_{2}$ edges, as all $2$-rainbow subgraphs containing $c_{0}$ or $c_{3}$ have size $0\mod 4$, but $\{c_{1},c_{2}\}$-rainbow subgraphs have size $6=2\mod 4$, and so $\{\cancel{c_{1}},c_{2}\}$- or $\{c_{1},\cancel{c_{2}}\}$-rainbows subgraphs would have size three. This fact can be understood as arising from the structure of the cubic lattice (which was used to obtain the original flag graph). This lattice has Coxeter diagram 434 which tells us that all $2$-rainbow subgraphs are generated by symmetries of even order except for $\{c_{1},c_{2}\}$-subgraphs which are generated by symmetries of odd order. This insight can allow us to perform edge contraction on colour codes defined on hyperbolic lattices, for example we demonstrate how to halve the number of physical qubits used in a hyperbolic colour code based on the 4-3-5 3D tiling in the linked Jupyter notebook. We can also see that, because the Coxeter diagram for a $D$-dimensional hypercubic lattice has the form 4334 we can in general only contract $c_{0}$ and $c_{D}$ edges in flag graphs obtained via the hypergraph product.
@@
This is not the only requirement, however. If we consider a graph $\mathcal{G}$ obtained from a product of a length-$4$ and a length-$6$ cycle which defines a [[48,4,4]][\![48,4,4]\!] 2D colour code then the graph $\mathrm{Cont}(c_{0}):\mathcal{G}$ defines a [[24,2,4]][\![24,2,4]\!] code, so two logical qubits have been lost. We can see this in Fig.˜15. We observe similar results in 3D, for example a product of three length-$6$ cycles defines a [[1296,9,6]][\![1296,9,6]\!] code but $c_{0}$ contraction results in a [[648,6,6]][\![648,6,6]\!] code and an additional $c_{3}$ contraction results in a [[324,6,6]][\![324,6,6]\!] code. This suggests that for the number of logical qubits to be preserved under contraction we should choose input graphs whose fundamental cycle bases contain only cycles of length $0\mod 4$ and not $2\mod 4$.
We also need to understand how contraction works at seams between colour codes. An example is shown in Fig.˜16. In this case we can see that, unlike in a colour code containing no seams, $\{\cancel{c_{0}},c_{1}\}$-rainbow subgraphs and $\{c_{1},c_{2}\}$-maximal subgraphs are not required to have even intersection and instead can intersect at a single vertex. This makes assignment of stabilisers to $c_{0}$-contracted lattices containing type-$1$ seams (or $c_{D}$-contracted lattices containing type-0 seams) quite difficult. In contrast, because $c_{D}$-contraction does not modify type-$1$ seams (and similarly for $c_{0}$ contraction and type-0 seams) the same stabiliser as in unjoined colour codes can be applied in this case. This means that in codes obtained from products of graphs in which all level-$1$ (level-0) vertices have degree $2$ we can reliably perform $c_{0}$ ($c_{D}$) edge contraction and reliably obtain a commuting stabiliser group.
Figure 16: (Left) Subgraph of the flag graph shown in Fig.˜6 that contains a seam. (Right) Transformation of this seam under $c_{0}$-contraction. The intersection of $\{\cancel{c_{0}},c_{1}\}$-maximal and $\{c_{1},c_{2}\}$-rainbow subgraphs can now be a single vertex.